# On the Provable Suboptimality of Momentum SGD in Nonstationary Stochastic Optimization

### v2

@@ 3.1 Bounds in expectation
    Run SGD with constant stepsize $\gamma_{k}$ for $T_{k}$ iterations in epoch $k$, starting from $\theta_{0}$. Let $T:=\sum_{k=0}^{K-1}T_{k}$ be the total horizon. Then the final iterate satisfies
@@
Theorem˜3.1 and Theorem˜3.2 give us algorithmic guarantees for running SGD in non-stationary environments. These closely match the guarantees provided by [CDH21] and are identical to the efficiency estimates in the static setting [LAN11]. Since the velocity iterate $\bm{v}_{t}$ is tightly coupled with $\bm{\psi}_{t}$, a direct adaptation of standard SGD-style arguments is not straightforward. One must instead analyze a coupled recursion involving both $\bm{v}_{t}$ and $\bm{\psi}_{t}$, which typically necessitates a Lyapunov (stability) function approach. However, constructing an appropriate Lyapunov function in this setting is itself cumbersome, making the resulting analysis considerably more delicate. In light of Lemma˜3.1, we can view SGD with momentum as a 2D dynamical system on extended state vectors. We defer the proof to Section˜A.2:
Lemma 3.1 (Extended 2D recursion for SGD with momentum).
@@
$\begin{bmatrix}\widehat{\bm{\theta}}_{t}\\ \check{\bm{\theta}}_{t}\end{bmatrix}=\begin{bmatrix}I_{d}-\frac{\gamma_{t}}{1-\beta}\mathbb{H}_{t-1}&\frac{\gamma_{t}\beta^{\prime}}{1-\beta}\mathbb{H}_{t-1}\\ -\frac{\gamma_{t}}{1-\beta}\mathbb{H}_{t-1}&\beta I_{d}+\frac{\gamma_{t}\beta^{\prime}}{1-\beta}\mathbb{H}_{t-1}\end{bmatrix}\begin{bmatrix}\widehat{\bm{\theta}}_{t-1}\\ \check{\bm{\theta}}_{t-1}\end{bmatrix}+\frac{1}{1-\beta}\begin{bmatrix}-(I_{d}-\gamma_{t}\mathbb{H}_{t-1})\Delta_{t-1}-\mathbb{K}_{t-1}\Delta_{t-2}\\ -(I_{d}-\gamma_{t}\mathbb{H}_{t-1})\Delta_{t-1}-\mathbb{K}_{t-1}\Delta_{t-2}\end{bmatrix}+\frac{\gamma_{t}}{1-\beta}\begin{bmatrix}\xi_{t+1}(\bm{\psi}_{t-1})\\ \xi_{t+1}(\bm{\psi}_{t-1})\end{bmatrix}$
@@
$\beta^{\prime}\stackrel{{\scriptstyle\Delta}}{{=}}\beta\beta_{1}+\beta_{2}$ (3.1)
@@
$\mathbb{H}_{t-1}\stackrel{{\scriptstyle\Delta}}{{=}}\int_{0}^{1}\nabla^{2}F_{t+1}(\bm{\theta}_{t+1}^{\star}+s(\bm{\psi}_{t}-\bm{\theta}_{t+1}^{\star}))ds$ (3.2)
$\mathbb{K}_{t-1}=-\beta I_{d}+\gamma_{t}\beta_{1}\mathbb{H}_{t-1}.$ (3.3)
This characterization of SGD with momentum proves very useful as it allows us to decouple the interaction between the noise induced from stochastic gradients, the error being accumulated by using momentum, and the effect of momentum on the drift. We now will provide an expectation bound on the tracking error for SGD with momentum. One will note that we incur extra scaling factors of $(1-\beta)^{-1}$ and $(1-\beta)^{-2}$ which explains why momentum can be significantly worse than SGD in non-stationary environments.
Theorem 3.3 (Tracking error bound in expectation for SGD with momentum).

### v3

@@ C.1 Proof for SGD tracking error bound
    Run (SGD) with constant stepsize $\gamma_{k}$ for $T_{k}$ iterations in epoch $k$, starting from $\theta_{0}$. Let $T:=\sum_{k=0}^{K-1}T_{k}$ be the total horizon. Then the final iterate satisfies
@@
Proof of Corollary˜C.2.
From Corollary˜C.1, we have the following
$\mathbb{E}\|\bm{\theta}_{t+1}-\bm{\theta}_{t+1}^{\star}\|^{2}\leq\left(1-\frac{\gamma\mu}{2}\right)^{t}\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}+\frac{4\Delta^{2}}{\gamma^{2}\mu^{2}}+\frac{\sigma^{2}\gamma}{\mu}.$ (7)
Proof of (i). By (7) and $\gamma=\gamma^{\star}$, using $\mathcal{E}(\gamma):=\sigma^{2}\gamma/\mu+4\Delta^{2}/\mu^{2}\gamma^{2}$, we have:
Proof of (ii). Let $t_{0}:=0$ and $t_{k+1}:=t_{k}+T_{k}$. Define epoch iterates $X_{k}:=\bm{\theta}_{t_{k}}$ and corresponding minimizers $X_{k}^{\star}:=\bm{\theta}_{t_{k}}^{\star}$. Applying (7) inside epoch $k$ with stepsize $\gamma_{k}$ and length $T_{k}$ yields
For $k\geq 1$, by the choice of $T_{k}$ and $\log(1-x)\leq-x$,
so for $k\geq 1$,
$\mathbb{E}\|X_{k+1}-X_{k+1}^{\star}\|^{2}\leq\frac{1}{4}\,\mathbb{E}\|X_{k}-X_{k}^{\star}\|^{2}+\mathcal{E}(\gamma_{k}).$ (8)
Induction. We claim that for all $k\geq 1$,
$\mathbb{E}\|X_{k}-X_{k}^{\star}\|^{2}\leq 2\mathcal{E}(\gamma_{k-1}).$
Assume this holds for some $k\geq 1$. Using (8),
$\mathcal{E}(2\gamma)\leq 2\mathcal{E}(\gamma).$
Next, by the definition of $K$,
Time bound. Since $\mathcal{E}(\gamma_{0})\geq\mathcal{E}$, we have
Moreover, for $k\geq 1$, $\gamma_{k}\geq\gamma_{0}/2^{k+1}$, hence
Thus
$\sum_{k=1}^{K-1}T_{k}\leq\sum_{k=1}^{K-1}\Bigl(\frac{2\log 4}{\mu\gamma_{k}}+1\Bigr)\leq\frac{8\log 4}{\mu}\sum_{k=1}^{K-1}\frac{1}{\gamma_{k}}+\mathcal{O}(K)\leq\frac{32\log 4}{\mu\gamma^{\star}}+\mathcal{O}(K).$
$\sum_{k=1}^{K-1}T_{k}\leq 32\log 4\cdot\frac{3}{2}\cdot\frac{\sigma^{2}}{\mu^{2}\mathcal{E}}+\mathcal{O}(1)=48\log 4\cdot\frac{\sigma^{2}}{\mu^{2}\mathcal{E}}+\mathcal{O}(1).$
C.2 Proof for SGDM tracking error bound
For (SGDM), we will take a different approach to prove bounds on the tracking error in expectation and with high-probability. We will instead work with a 2D state-space view of (SGDM) by defining extended state vectors. We will first introduce the transformation matrices:
$\mathbb{V}=\begin{bmatrix}I_{d}&-\beta I_{d}\\ I_{d}&-I_{d}\end{bmatrix},\;\mathbb{V}^{-1}=\frac{1}{1-\beta}\begin{bmatrix}I_{d}&-\beta I_{d}\\ I_{d}&-I_{d}\end{bmatrix}$. (9)
Recall the following (SGDM) updates:
$\begin{split}&\bm{\psi}_{t}=\bm{\theta}_{t}+\beta_{1}(\bm{\theta}_{t}-\bm{\theta}_{t-1})\\ &\bm{\theta}_{t+1}=\bm{\psi}_{t}-\gamma_{t}m_{t+1}(\bm{\psi}_{t})-\gamma_{t}\bm{\xi}_{t+1}(\bm{\psi}_{t})+\beta_{2}(\bm{\psi}_{t}-\bm{\psi}_{t-1}).\end{split}$ (10)
$\begin{bmatrix}\widehat{\bm{\theta}}_{t}\\ \check{\bm{\theta}}_{t}\end{bmatrix}\stackrel{{\scriptstyle\Delta}}{{=}}\mathbb{V}^{-1}\begin{bmatrix}\widetilde{\bm{\theta}}_{t}\\ \widetilde{\bm{\theta}}_{t-1}\end{bmatrix}=\frac{1}{1-\beta}\begin{bmatrix}\widetilde{\bm{\theta}}_{t}-\beta\widetilde{\bm{\theta}}_{t-1}\\ \widetilde{\bm{\theta}}_{t}-\widetilde{\bm{\theta}}_{t-1}\end{bmatrix}$. (11)
We can obtain an extended recursion 2D state-space matrix that captures the dynamics of (SGDM). This will prove very useful in our subsequent analysis to obtain expectation and high probability bounds without needing to recurse on the velocity vector $\mathbb{v}_{t}$ or using a Lyapunov stability function argument. Before we proceed with obtaining this extended recursion, we state an assumption that typically holds under sufficient conditions for the Dominated convergence theorem to hold:
Assumption C.2 (Interchanging conditional expectation and gradient).
For each $t\geq 0$, define the conditional objective
Assume $F_{t+1}$ is differentiable and
Note that since we assumed $m_{t+1}$ is $\mu$-strongly monotone and Lipschitz continuous (˜B.2 and ˜B.3), it follows that $F_{t+1}$ is $\mu$-strongly convex and $L$-smooth. Equivalently, this implies
$\mu I_{d}\preceq\nabla^{2}F_{t+1}(\bm{\theta})\preceq LI_{d}\;\;\;\text{$\forall\bm{\theta}\in\mathbb{R}^{d}$}.$ (12)
We now prove the extended 2D recursion. This proof largely follows (Yuan et al., 2016) with a similar form except for a matrix decoupling that arises from the minimizer drift.
Lemma C.2 (Extended 2D recursion for SGD with momentum).
@@ C.2 Proof for SGDM tracking error bound
$\begin{bmatrix}\widehat{\bm{\theta}}_{t}\\ \check{\bm{\theta}}_{t}\end{bmatrix}=\begin{bmatrix}I_{d}-\frac{\gamma_{t}}{1-\beta}\mathbb{H}_{t-1}&\frac{\gamma_{t}\beta^{\prime}}{1-\beta}\mathbb{H}_{t-1}\\ -\frac{\gamma_{t}}{1-\beta}\mathbb{H}_{t-1}&\beta I_{d}+\frac{\gamma_{t}\beta^{\prime}}{1-\beta}\mathbb{H}_{t-1}\end{bmatrix}\begin{bmatrix}\widehat{\bm{\theta}}_{t-1}\\ \check{\bm{\theta}}_{t-1}\end{bmatrix}+\frac{1}{1-\beta}\begin{bmatrix}-(I_{d}-\gamma_{t}\mathbb{H}_{t-1})\Delta_{t-1}-\mathbb{K}_{t-1}\Delta_{t-2}\\ -(I_{d}-\gamma_{t}\mathbb{H}_{t-1})\Delta_{t-1}-\mathbb{K}_{t-1}\Delta_{t-2}\end{bmatrix}+\frac{\gamma_{t}}{1-\beta}\begin{bmatrix}\bm{\xi}_{t}(\bm{\psi}_{t-1})\\ \bm{\xi}_{t}(\bm{\psi}_{t-1})\end{bmatrix}$ (13)
@@
$\beta^{\prime}\stackrel{{\scriptstyle\Delta}}{{=}}\beta\beta_{1}+\beta_{2}$ (14)
$\mathbb{H}_{t-1}\stackrel{{\scriptstyle\Delta}}{{=}}\int_{0}^{1}\nabla^{2}F_{t+1}(\bm{\theta}_{t+1}^{\star}+s(\bm{\psi}_{t}-\bm{\theta}_{t+1}^{\star}))ds$ (15)
$\mathbb{K}_{t-1}=-\beta I_{d}+\gamma_{t}\beta_{1}\mathbb{H}_{t-1}.$ (16)
Proof of Lemma˜C.2.
From the (SGDM) update, we have
$\bm{\theta}_{t+1}=\bm{\psi}_{t}-\gamma_{t}m_{t+1}(\bm{\psi}_{t})-\gamma_{t}\bm{\xi}_{t+1}(\bm{\psi}_{t})+\beta_{2}(\bm{\psi}_{t}-\bm{\psi}_{t-1}).$ (17)
Let $\widetilde{\bm{\theta}}_{t}=\bm{\theta}_{t}^{\star}-\bm{\theta}_{t}$ and $\widetilde{\bm{\psi}}_{t}=\bm{\theta}_{t+1}^{\star}-\bm{\psi}_{t}$. Subtracting both sides of (17) from $\bm{\theta}_{t}^{\star}$, we get:
$\widetilde{\bm{\theta}}_{t+1}=\widetilde{\bm{\psi}}_{t}+\gamma_{t}m_{t+1}(\bm{\psi}_{t})+\gamma_{t}\bm{\xi}_{t+1}(\bm{\psi}_{t})-\beta_{2}(\bm{\psi}_{t}-\bm{\psi}_{t-1}).$ (18)
We can now appeal to the mean-value theorem:
@@
$\nabla F_{t+1}(\bm{\psi}_{t})-\nabla F_{t+1}(\bm{\theta}_{t+1}^{\star})=\left(\int_{0}^{1}\nabla^{2}F_{t+1}(\bm{\theta}_{t+1}^{\star}+s(\bm{\psi}_{t}-\bm{\theta}_{t+1}^{\star}))ds\right)(\bm{\psi}_{t}-\bm{\theta}_{t+1}^{\star})\stackrel{{\scriptstyle\Delta}}{{=}}-\mathbb{H}_{t}\widetilde{\bm{\psi}}_{t}.$ (19)
$\bm{\psi}_{t}-\bm{\psi}_{t-1}=(\widetilde{\bm{\psi}}_{t}-\bm{\theta}_{t+1}^{\star})-(\widetilde{\bm{\psi}}_{t-1}-\bm{\theta}_{t+1}^{\star})=-\widetilde{\bm{\psi}}_{t}+\widetilde{\bm{\psi}}_{t-1}.$ (20)
Combining (18), (20) and (19), we find
$\begin{split}\widetilde{\bm{\theta}}_{t+1}&=\widetilde{\bm{\psi}}_{t}+\gamma_{t}m_{t+1}(\bm{\psi}_{t})+\gamma_{t}\bm{\xi}_{t+1}(\bm{\psi}_{t})-\beta_{2}(\bm{\psi}_{t}-\bm{\psi}_{t-1})\\ &=\widetilde{\bm{\psi}}_{t}-\gamma_{t}\mathbb{H}_{t}\widetilde{\bm{\psi}}_{t}+\beta_{2}(\widetilde{\bm{\psi}}_{t}-\widetilde{\bm{\psi}}_{t-1})+\gamma_{t}\bm{\xi}_{t+1}(\bm{\psi}_{t})\\ &=((1+\beta_{2})I_{d}-\gamma_{t}\mathbb{H}_{t})\widetilde{\bm{\psi}}_{t}-\beta_{2}\widetilde{\bm{\psi}}_{t-1}+\gamma_{t}\bm{\xi}_{t+1}(\bm{\psi}_{t}).\end{split}$ (21)
Now for any $k\leq t$, let us define
$\widetilde{\bm{\theta}}_{k}^{(t+1)}:=\bm{\theta}_{t+1}^{\star}-\bm{\theta}_{k}.$ (22)
Then we have that $\widetilde{\bm{\theta}}_{t+1}^{(t+1)}=\widetilde{\bm{\theta}}_{t+1}$. Using the lookahead from (SGDM), we have
$\widetilde{\bm{\psi}}_{t}=\bm{\theta}_{t+1}^{\star}-\bm{\psi}_{t}=(\bm{\theta}_{t+1}^{\star}-\bm{\theta}_{t})-\beta_{1}\left(\bm{\theta}_{t}-\bm{\theta}_{t-1}\right)=\widetilde{\bm{\theta}}_{t}^{(t+1)}-\beta_{1}\left(\bm{\theta}_{t}-\bm{\theta}_{t-1}\right).$ (23)
However also note that
$\bm{\theta}_{t}-\bm{\theta}_{t-1}=-(\bm{\theta}_{t+1}^{\star}-\bm{\theta}_{t})+(\bm{\theta}_{t+1}^{\star}-\bm{\theta}_{t-1})=-\widetilde{\bm{\theta}}_{t}^{(t+1)}+\widetilde{\bm{\theta}}_{t-1}^{(t+1)}.$ (24)
Combining (23) and (24), we find
$\widetilde{\bm{\psi}}_{t}=\widetilde{\bm{\theta}}_{t}^{(t+1)}-\beta_{1}\left(-\widetilde{\bm{\theta}}_{t}^{(t+1)}+\widetilde{\bm{\theta}}_{t-1}^{(t+1)}\right)=(1+\beta_{1})\widetilde{\bm{\theta}}_{t}^{(t+1)}-\beta_{1}\widetilde{\bm{\theta}}_{t-1}^{(t+1)}.$ (25)
Plugging in (25) into (21), we obtain the following:
$\widetilde{\bm{\theta}}_{t+1}^{(t+1)}=\mathbb{J}_{t}\widetilde{\bm{\theta}}_{t}^{(t+1)}+\mathbb{K}_{t}\widetilde{\bm{\theta}}_{t-1}^{(t+1)}+L\widetilde{\bm{\theta}}_{t-2}^{(t+1)}+\gamma_{t}\bm{\xi}_{t+1}(\bm{\psi}_{t}),$ (26)
where
$\begin{split}&\mathbb{J}_{t}\stackrel{{\scriptstyle\Delta}}{{=}}(1+\beta_{1})(1+\beta_{2})I_{d}-\gamma_{t}(1+\beta_{1})\mathbb{H}_{t}\\ &\mathbb{K}_{t}\stackrel{{\scriptstyle\Delta}}{{=}}-(\beta_{1}+\beta_{2}+\beta_{1}\beta_{2})I_{d}+\gamma_{t}\beta_{1}\mathbb{H}_{t}\\ &L\stackrel{{\scriptstyle\Delta}}{{=}}\beta_{1}\beta_{2}.\end{split}$ (27)
$\begin{split}&\mathbb{J}_{t}=(1+\beta)I_{d}-\gamma_{t}(1+\beta_{1})\mathbb{H}_{t}\\ &\mathbb{K}_{t}=-\beta I_{d}+\gamma_{t}\beta_{1}\mathbb{H}_{t}\\ &L=0.\end{split}$ (28)
$\widetilde{\bm{\theta}}_{t}^{(t+1)}=\bm{\theta}_{t+1}^{\star}-\bm{\theta}_{t}=(\bm{\theta}_{t}^{\star}-\bm{\theta}_{t})+(\bm{\theta}_{t+1}^{\star}-\bm{\theta}_{t}^{\star})=\widetilde{\bm{\theta}}_{t}-\Delta_{t}.$ (29)
Similarly we have
$\widetilde{\bm{\theta}}_{t-1}^{(t+1)}=(\bm{\theta}_{t-1}^{\star}-\bm{\theta}_{t-1})+(\bm{\theta}_{t+1}^{\star}-\bm{\theta}_{t}^{\star}+\bm{\theta}_{t}^{\star}-\bm{\theta}_{t-1}^{\star})=\widetilde{\bm{\theta}}_{t-1}-(\Delta_{t-1}+\Delta_{t}).$ (30)
Plugging (29) and (30) into (26), we get
$\widetilde{\bm{\theta}}_{t+1}=\mathbb{J}_{t}\widetilde{\bm{\theta}}_{t}+\mathbb{K}_{t}\widetilde{\bm{\theta}}_{t-1}+\mathbb{b}_{t}+\gamma_{t}\bm{\xi}_{t+1}(\bm{\psi}_{t}),$ (31)
where
$\mathbb{b}_{t}\stackrel{{\scriptstyle\Delta}}{{=}}-(\bm{J}_{t}+\bm{K}_{t})\Delta_{t}-\bm{K}_{t}\Delta_{t-1}=-(I_{d}-\gamma_{t}\mathbb{}H_{t})\Delta_{t}-\mathbb{K}_{t}\Delta_{t-1}.$ (32)
It follows that we can write the extended recursion relation as:
$\underbrace{\begin{bmatrix}\widetilde{\bm{\theta}}_{t+1}\\ \widetilde{\bm{\theta}}_{t}\end{bmatrix}}_{\stackrel{{\scriptstyle\Delta}}{{=}}\;\mathbb{z}_{t+1}}=\underbrace{\begin{bmatrix}\mathbb{J}_{t}&\mathbb{K}_{t}\\ I_{d}&0\end{bmatrix}}_{\stackrel{{\scriptstyle\Delta}}{{=}}\;\mathbb{B}_{t}}\underbrace{\begin{bmatrix}\widetilde{\bm{\theta}}_{t}\\ \widetilde{\bm{\theta}}_{t-1}\end{bmatrix}}_{{\stackrel{{\scriptstyle\Delta}}{{=}}\;\mathbb{z}_{t}}}+\underbrace{\begin{bmatrix}\mathbb{b}_{t}\\ 0\end{bmatrix}}_{\stackrel{{\scriptstyle\Delta}}{{=}}\;\mathbb{u}_{t}}+\gamma_{t}\underbrace{\begin{bmatrix}\bm{\xi}_{t+1}(\bm{\psi}_{t})\\ 0\end{bmatrix}}_{\stackrel{{\scriptstyle\Delta}}{{=}}\;\bm{\eta}_{t+1}}.$ (33)
Note that we can write $\mathbb{B}_{t}=\mathbb{P}-\mathbb{M}_{t}$ where
$\mathbb{P}=\begin{bmatrix}(1+\beta)I_{d}&-\beta I_{d}\\ I_{d}&0\end{bmatrix},\;\;\mathbb{M}_{t}=\begin{bmatrix}\gamma_{t}(1+\beta_{1})\mathbb{H}_{t}&-\gamma_{t}\beta_{1}\mathbb{H}_{t}\\ 0&0\end{bmatrix}$. (34)
$\mathbb{V}=\begin{bmatrix}I_{d}&-\beta I_{d}\\ I_{d}&-I_{d}\end{bmatrix},\;\;\mathbb{V}^{-1}=\frac{1}{1-\beta}\begin{bmatrix}I_{d}&-\beta I_{d}\\ I_{d}&-I_{d}\end{bmatrix},\;\;\mathbb{D}=\begin{bmatrix}I_{d}&0\\ 0&\beta I_{d}\end{bmatrix}$. (35)
Define the transformed "mode-splitting" coordinates as follows:
$\mathbb{y}_{t}:=\mathbb{V}^{-1}\mathbb{z}_{t}=\begin{bmatrix}\widehat{\bm{\theta}}_{t}\\ \check{\bm{\theta}}_{t}\end{bmatrix}=\frac{1}{1-\beta}\begin{bmatrix}\widetilde{\bm{\theta}}_{t}-\beta\widetilde{\bm{\theta}}_{t-1}\\ \widetilde{\bm{\theta}}_{t}-\widetilde{\bm{\theta}}_{t-1}\end{bmatrix}$. (36)
Using this transform with (31) gives us:
$\begin{split}\mathbb{y}_{t+1}&=\mathbb{V}^{-1}\mathbb{B}_{t}\mathbb{V}\mathbb{y}_{t}+\mathbb{V}^{-1}\mathbb{u}_{t}+\gamma_{t}\mathbb{V}^{-1}\bm{\eta}_{t+1}\\ &=\underbrace{(\mathbb{D}-\mathbb{V}^{-1}\mathbb{M}_{t}\mathbb{V})}_{\stackrel{{\scriptstyle\Delta}}{{=}}\;\widetilde{\mathbb{B}}_{t}}\mathbb{y}_{t}+\mathbb{V}^{-1}\mathbb{u}_{t}+\gamma_{t}\mathbb{V}^{-1}\bm{\eta}_{t+1}.\end{split}$ (37)
Finally we note that
$\mathbb{D}-\mathbb{V}^{-1}\mathbb{M}_{t}\mathbb{V}=\begin{bmatrix}I_{d}-\frac{\gamma_{t}}{1-\beta}\mathbb{H}_{t}&\frac{\gamma_{t}\beta^{\prime}}{1-\beta}\mathbb{H}_{t}\\ -\frac{\gamma_{t}}{1-\beta}\mathbb{H}_{t}&\beta I_{d}+\frac{\gamma_{t}\beta^{\prime}}{1-\beta}\mathbb{H}_{t}\end{bmatrix},\;\;\mathbb{V}^{-1}\bm{\eta}_{t+1}=\frac{1}{1-\beta}\begin{bmatrix}\bm{\xi}_{t+1}(\bm{\psi}_{t})\\ \bm{\xi}_{t+1}(\bm{\psi}_{t})\end{bmatrix}$. (38)
This concludes the proof. ∎
In order to establish a convergence guarantee of (SGDM), we will need to analyze the dynamics of the norm of the extended state vectors. What we aim to do is show a path-wise inequality (analogous to a Lyapunov stability function) of the form $\mathbb{s}_{t}\leq\bm{\Gamma}_{t}\mathbb{s}_{t}$ and then get a uniform $\bm{\Gamma}$ by upper bounding $\bm{\Gamma}_{t}$, $\forall t\geq 0$.
Corollary C.3 (Uniform stability for (SGDM)).